How many dimensions does hillbert space of 10 identical particles?
1
Expert's answer
2020-06-17T09:38:39-0400

A Hilbert space is a vector space with an inner product such that the norm defined by "\begin{vmatrix}
 f\\
\end{vmatrix}" = square root of (f, f) turns H into a complete metric space. If the metric defined by the norm is not complete, then H is known as an inner product space.

Examples of finite-dimensional Hilbert spaces include:

Real numbers Rn with (v, u) the vector dot product of "\nu" and u

Complex numbers Cn with (v, u), the vector dot product of "\nu" and u and the complex conjugate of u.

As a consequence of Zorn's lemma, every Hilbert space admits an orthonormal basis; any two orthonormal bases of the same space have the same cardinality, called Hilbert dimension of the space. For example, I2 (B)

has an orthonormal basis indexed by B, its Hilbert dimension is the cardinality of B (which may be a finite integer or a countable or uncountable cardinal number).

As a consequence of Parseval's identity, if {ek}k ∈ B is an orthonormal basis of H, then the map Φ : H →l2(B) ( defined by Φ(x) = ⟨x, ek⟩k∈B is an isometric isomorphism of Hilbert spaces: it is a bijective linear mapping such that

al basis indexed by B, its Hilbert dimension is the cardinality of B (which may be a finite integer, or a countable or uncountable cardinal number).

As a consequence of Parseval's identity, if {ek}k ∈ B is an orthonormal basis of H, then the map Φ : H → l2(B) defined by Φ(x) = ⟨x, ek⟩k∈B is an isometric isomorphism of Hilbert spaces: it is a bijective linear mapping such that"{\displaystyle {\bigl \langle }\Phi (x),\Phi (y){\bigr \rangle }_{l^{2}(B)}=\left\langle x,y\right\rangle _{H}}"

for all x, y ∈ H. The cardinal number of B is the Hilbert dimension of H. So, every Hilbert space is isometrically isomorphic to a sequence space  l2(B) for some set B.

The sequence space l2 consists of all infinite sequences z = (z1, z2, …) of complex numbers such that the series


"{\displaystyle \sum _{n=1}^{\infty }|z_{n}|^{2}}"

converges.

The inner product on I^2 is defined by

"{\displaystyle \langle \mathbf {z} ,\mathbf {w} \rangle =\sum _{n=1}^{\infty }z_{n}{\overline {w_{n}}}\,,}"

with the latter series converging as a consequence of Cauchy-Schwarz inequality.

Completeness of the space holds provided that whenever a series of elements from I^2 converges absolutely, then it converges to an element of I^2.

Consider two particles occupying N distinct states |1>, 2>,....,| N>. The Hilbert space dimension of this two-particle system if the two particles are identical bosons will be just N and N^2 - N due to Pauli exclusion if they are in different states.

Assume the states to be N*1column matrices. The two particle state is an N*N

matrix. For bosons, we need a symmetrical matrix. It has N(N+1)/2 distinct terms. For fermions, we need an anti-symmetrical matrix that has N(N-1)/2 distinct terms.